# OmniCompliance-100K: A Multi-Domain, Rule-Grounded, Real-World Safety Compliance Dataset

### 14 Mar 2026 v1

@@ Abstract
Ensuring the safety and compliance of large language models (LLMs) is of paramount importance. However, existing LLM safety datasets often rely on ad-hoc taxonomies for data generation and suffer from a significant shortage of rule-grounded, real-world cases that are essential for robustly protecting LLMs. In this work, we address this critical gap by constructing a comprehensive safety dataset from a compliance perspective. Using a powerful web-searching agent, we collect a rule-grounded, real-world case dataset OmniCompliance-100K, sourced from multi-domain authoritative references. The dataset spans 74 regulations and policies across a wide range of domains, including security and privacy regulations, content safety and user data privacy policies from leading AI companies and social media platforms, financial security requirements, medical device risk management standards, educational integrity guidelines, and protections of fundamental human rights. In total, our dataset contains 12,985 distinct rules and 106,009 associated real-world compliance cases. Our analysis confirms a strong alignment between the rules and their corresponding cases. We further conduct extensive benchmarking experiments to evaluate the safety and compliance capabilities of advanced LLMs across different model scales. Our experiments reveal several interesting findings that have great potential to offer valuable insights for future LLM safety research.
@@
Wenbin Huthanks: Equal Contribution, Huihao Jing, Haochen Shi, Changxuan Fan, Haoran Lithanks: Corresponding author, Yangqiu Song Hong Kong University of Science and Technology, {whuak,hjingaa,hshiah,cfanam,hlibt}@connect.ust.hk, yqsong@cse.ust.hk
@@ 1 Introduction
With the rapid deployment of large language models (LLMs) DeepSeek-AI et al. (2025a); Ouyang et al. (2022); Touvron et al. (2023), across diverse industries, their potential risks have become increasingly prominent. From generating harmful content Liu et al. (2024) and leaking private information Kim et al. (2023), to violating financial compliance requirements Chen et al. (2025), misleading medical decisions Alber et al. (2025), or infringing upon fundamental human rights Raman et al. (2025), LLMs can easily cause serious social, economic, and legal consequences in real-world scenarios. Ensuring the safety of LLMs has thus emerged as one of the most urgent issues in both academia and industry.
@@
Meanwhile, researchers are working to address safety issues by ensuring that LLMs comply with established AI safety regulations and policies, as they are carefully designed by legal experts and authoritative organizations, providing comprehensive guidelines that address a wide range of risks. Air-Bench Zeng et al. (2024) has developed a safety taxonomy based on government regulations and company policies, which it then uses to synthesize a safety dataset with LLMs. Another work, GuardSet-X Kang et al. (2025), gathers policies from safety-sensitive domains and also synthesizes cases based on these policies using LLMs. While these works represent innovative steps toward addressing safety issues based on established safety rules, they are notably deficient in real-world cases. Synthesized data leads to a lack of diversity and poor generalization in real-world applications.
@@ 2.1 Safety Datasets
Researchers have made efforts to create safety datasets for evaluating or aligning LLMs. However, these datasets are usually generated by LLMs based on ad-hoc safety taxonomies Ji et al. (2023); Mazeika et al. (2024); Ricciuti (2025). For example, ToxicChat Lin et al. (2023) is produced by Vicuna Chiang et al. (2023), and WildGuard Han et al. (2024) is created by GPT-4 OpenAI et al. (2024b), making their generalization in real-world applications challenging. Furthermore, relying on ad-hoc safety taxonomies leads to a lack of systematic and rigorous protection for LLM safety.
@@ 2.2 Compliance Datasets
On the other hand, as new regulations and policies regarding LLM safety are developed, these sources provide invaluable expert compliance guidelines for ensuring safe LLM usage. Recently, researchers have emphasized the importance of addressing safety from a compliance perspective Hu et al. (2025). However, existing datasets are either limited in scale or lack real compliance cases. AirBench Zeng et al. (2024) creates a safety taxonomy based on regulations and generates more than 5,000 cases accordingly, while GuardSet-X Kang et al. (2025) directly creates more than 120,000 cases generated based on their collected rules. Unfortunately, both methods fail to capture real-world cases, which restricts the generalizability of their datasets. Another effort, PrivaCI-Bench Li et al. (2025), contains around 3,000 real court cases. However, it is not specifically related to LLM safety, and the cases are not well-grounded with the rules. In this work, we propose a real-world, rule-grounded, large-scale safety compliance dataset, called OmniCompliance-100K. As illustrated in Table 1, we compare these existing datasets with ours along several dimensions, including the number of rules, the number of cases, and whether they contain real cases.
@@ 4.4 Article Correlation in GDPR
In this section, we assess the correlations among GDPR articles by utilizing the rule-case knowledge graph (KG), as referenced in Section 3.3. We present these article correlations in a confusion matrix, as illustrated in Figure 5. This matrix captures all triplets of <article, searched case, reference article>. We then normalize the confusion matrix on a logarithmic scale ranging from 0 to 1.

### 16 Apr 2026 v2

@@ Abstract
Ensuring the safety and compliance of large language models (LLMs) is of paramount importance. However, existing LLM safety datasets often rely on ad-hoc taxonomies for data generation and suffer from a significant shortage of rule-grounded, real-world cases that are essential for robustly protecting LLMs. In this work, we address this critical gap by constructing a comprehensive safety dataset from a compliance perspective. Using a powerful web-searching agent, we collect a rule-grounded, real-world case dataset OmniCompliance-100K, sourced from multi-domain authoritative references. The dataset spans 74 regulations and policies across a wide range of domains, including security and privacy regulations, content safety and user data privacy policies from leading AI companies and social media platforms, financial security requirements, medical device risk management standards, educational integrity guidelines, and protections of fundamental human rights. In total, our dataset contains 12,985 distinct rules and 106,009 associated real-world compliance cases. Our analysis confirms a strong alignment between the rules and their corresponding cases. We further conduct extensive benchmarking experiments to evaluate the safety and compliance capabilities of advanced LLMs across different model scales. Our experiments reveal several interesting findings that have great potential to offer valuable insights for future LLM safety research. Our source code are available at https://github.com/HKUST-KnowComp/OmniCompliance-100K.
@@
Wenbin Hu1,2thanks: Equal Contribution, Huihao Jing1,2∗, Haochen Shi1, Changxuan Fan1, Haoran Li1thanks: Corresponding author, Yangqiu Song1,2 1Hong Kong University of Science and Technology, 2ModeIO {whuak,hjingaa,hshiah,cfanam,hlibt}@connect.ust.hk, yqsong@cse.ust.hk
@@ 1 Introduction
With the rapid deployment of large language models (LLMs) DeepSeek-AI et al. (2025a); Touvron et al. (2023), across diverse industries, their potential risks have become increasingly prominent. From generating harmful content Liu et al. (2024) and leaking private information Kim et al. (2023), to violating financial compliance requirements Chen et al. (2025), misleading medical decisions Alber et al. (2025), or infringing upon fundamental human rights Raman et al. (2025), LLMs can easily cause serious social, economic, and legal consequences in real-world scenarios. Ensuring the safety of LLMs has thus emerged as one of the most urgent issues in both academia and industry.
@@
Meanwhile, researchers are working to address safety issues by ensuring that LLMs comply with established AI safety regulations and policiesHu et al. (2025b, a), as they are carefully designed by legal experts and authoritative organizations, providing comprehensive guidelines that address a wide range of risks. Air-Bench Zeng et al. (2024) has developed a safety taxonomy based on government regulations and company policies, which it then uses to synthesize a safety dataset with LLMs. Another work, GuardSet-X Kang et al. (2025), gathers policies from safety-sensitive domains and also synthesizes cases based on these policies using LLMs. While these works represent innovative steps toward addressing safety issues based on established safety rules, they are notably deficient in real-world cases. Synthesized data leads to a lack of diversity and poor generalization in real-world applications.
@@ 2.1 Safety Datasets
Researchers have made efforts to create safety datasets for evaluating or aligning LLMs. However, these datasets are usually generated by LLMs based on ad-hoc safety taxonomies Mazeika et al. (2024); Jing et al. (2025). For example, ToxicChat Lin et al. (2023) is produced by Vicuna Chiang et al. (2023), and WildGuard Han et al. (2024) is created by GPT-4 OpenAI et al. (2024b), making their generalization in real-world applications challenging. Furthermore, relying on ad-hoc safety taxonomies leads to a lack of systematic and rigorous protection for LLM safety.
@@ 2.2 Compliance Datasets
On the other hand, as new regulations and policies regarding LLM safety are developed, these sources provide invaluable expert compliance guidelines for ensuring safe LLM usage. Recently, researchers have emphasized the importance of addressing safety from a compliance perspective Hu et al. (2025a). However, existing datasets are either limited in scale or lack real compliance cases. AirBench Zeng et al. (2024) creates a safety taxonomy based on regulations and generates more than 5,000 cases accordingly, while GuardSet-X Kang et al. (2025) directly creates more than 120,000 cases generated based on their collected rules. Unfortunately, both methods fail to capture real-world cases, which restricts the generalizability of their datasets. Another effort, PrivaCI-Bench Li et al. (2025), contains around 3,000 real court cases. However, it is not specifically related to LLM safety, and the cases are not well-grounded with the rules. In this work, we propose a real-world, rule-grounded, large-scale safety compliance dataset, called OmniCompliance-100K. As illustrated in Table 1, we compare these existing datasets with ours along several dimensions, including the number of rules, the number of cases, and whether they contain real cases.
@@ 4.4 Article Correlation in GDPR
In this section, we assess the correlations among GDPR articles by utilizing the rule-case knowledge graph (KG), as referenced in Section 3.3. We present these article correlations in a confusion matrix $M$, as illustrated in Figure 5. This matrix captures all triplets of <article, searched case, reference article>. We then normalize the confusion matrix on a logarithmic scale ranging from 0 to 1, by calculating $\frac{\log(M_{ij}+1)}{\log(M_{\max}+1)}$, where $M_{\max}$ is the maximum value in the matrix $M$.
@@ 6 Conclusion
Acknowledgments
The authors of this paper were supported by the National Key Research and Development Program of China (2025YFE0200500), the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong, SAR, China.